Which of the following pairs represent the same rational number?
(8/-5) and (-24/15)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to check the given pair represents the same rational number.
​Then,
8/-5 = -24/15
8/-5 = -8/5
∵-8/5 = -8/5
∴8/-5 = -24/15
So, the given pair is represents the same rational number.

​​Thanks